OpenStudy (anonymous):

can you help me with this limit when x tends to zero, (1/x)(1/√(1+x^-2)), I find an indeterminate form for the denominator

OpenStudy (across):

\[\lim_{x\to0}\frac1x\frac1{\sqrt{1+x^{-2}}}?\]

OpenStudy (across):

If so, after performing some algebra, you see that this is the same as\[\lim_{x\to0}\frac{1}{\sqrt{x^2+1}}.\]

OpenStudy (anonymous):

it becomes \[1/\sqrt{1+x^2}\]

OpenStudy (anonymous):

put x=0 and u have ur answer

OpenStudy (anonymous):

the limit is 1 thanks guys!